OpenStudy (anonymous):

Solve. The icon indicates an exercise designed to give practice using a calculator.{3/4}x=18

OpenStudy (mertsj):

Multiply both sides by 4/3

OpenStudy (anonymous):

x = 24

OpenStudy (anonymous):

Thanks

jhonyy9 (jhonyy9):

x=18(4/3) x=6*4 x=24